I have to design an insulated tank which can keep the grains for 65oC for 1 hour. outside temp 25oC, volume of the tank 5000L cp=4.1 kj/kgK, p=1060kg/m3. optimize the height and radius of the tank.
Answer #1
Inputs
1) Inside Temp -Tin = 65 C
2) Outside Temp - T out = 25 C
3) Volume of tank - V= 5000 lit
4) cp =4.1 Kj/kg K
5) density - p = 1060 kg/m^3
Solution
In this problem,
ther are two constaints
a) volume of tank - 5000 lit = 5 m3
b) for minimum heat loss - surface area should be minimum
If r is radius of tank and h is height ,
volume V is given by
V = \tiny \prod r2h
And
surface area is given by
A = 2 r2 + 2 rh
Here A is objective that is to be minimized and V = 5 m^3
represents the constraint
From constaint equation h = V/( r2)
substituting this relationship into expression for Area will give
A = 2 ∏ r2 + 2∏ r V/( ∏ r2)
A = 2 ∏ r2 + 2 V/ r
Differentaiting A with respect to r and setting derivative equal to zero
to obtain the radius to optimum value ,
We get
dA/ dr = 4 ∏ r - 2 V/ r2 = 0
4 ∏ r = 2 V/ r2
4 ∏ r3 = 2 V
r3 = V /2∏
r = ( V /2∏)1/3 &
h =V/ ∏r2 = (4V/∏ )1/3
Putting V =5 m3
r = (5/2 x ∏)1/3 = 0.9266 m
h = (4V/∏ )1/3 =(4*5/∏)1/3 = 1.853 m
A = 2 ∏ r2 + 2∏ rh =16.189 m2
Second derivative is calculated to determine nature of the optimum.
Thus
\tiny \frac{\partial^2 }{\partial x^2} A = 4 x ∏ + 4 V/r3
As r is positive, second derivative is also positive, it indicates that surface area is a
minimum as required for mimimum heat loss.
Thus minimum Radius - r = 0.9266 m
and minimum height - h = 1.853 m
for maintaing condition of minimum heat loss that grain temp is kept at 65 C for 1 hour